The length of a rectangle is 16 feet more than twice its width. If the perimeter is 104, find the length of the rectangle.
Answer by checkley77(12844)

L=2W+16
2L+2W=104
2(2W+16)+2W=104
4W+32+2W=104
6W=104-32
6W=72
W=72/6
W=12 ANS. FOR THE WIDTH.
L=2*12+16=24+16=40 ANS FOR THE LENGTH.
PROOF:
2*40+2*12=104
80+24=104
104=104